OpenStudy (anonymous):

Match each equation with the ordered pair that is a solution. Match Term Definition y equals x plus two A) five comma four y equals two x minus six B) negative two comma two y equals three x plus eight C) negative six comma negative four y equals x minus seven D) four comma negative three

Directrix (directrix):

Use mathematical symbols. y equals x plus two A) five comma four y = x + 2 A) (5,4) is easier on the eyes. Agree? @nyprincess2001

OpenStudy (anonymous):

y=x+2 A.) (5,4) y=2x-6 B.) (-2,2) y=3x+8 C.) (-6,-4) y=x-7 D.) (4,-3)

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Start off with this: y=x+2 A.) (5,4) In y = x + 2, replace x by 5 and y by 4 and see if a true statement results. Does 4 = 5 + 2? Yes or No? No, because 4 is not equal to 7.

Directrix (directrix):

Take the same equation: y=x+2 and test it with the next point B.) (-2,2) Does 2 = -2 + 2 Yes or No? No, because 2 is not equal to 0.

OpenStudy (anonymous):

Oh okay you could take a break from me I got this question lol

Directrix (directrix):

Take the same equation: y=x+2 and test it with the next point C.) (-6,-4) Does -4 = -6 + 2 ? Yes or No. YES, because -4 = -4.

Directrix (directrix):

So, this --> y=x+2 goes with C. @nyprincess2001

OpenStudy (anonymous):

Yeah, thank You

Directrix (directrix):

Your turn: y=2x-6 Check it for A.) (5,4)

Directrix (directrix):

Oh, I just saw this: Oh okay you could take a break from me I got this question lol Am I getting on your nerves?

OpenStudy (anonymous):

no i feel like im getting on your nerves by asking you all these question lol

Directrix (directrix):

No,not at all. Go ahead and do this one: y=2x-6 Check it for A.) (5,4)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I GOT EVERYTHING THANK YOU

Directrix (directrix):

You are welcome.